### v1

# Predicting User Behavior in Smart Spaces with LLM-Enhanced Logs and Personalized Prompts

Abstract
Enhancing the intelligence of smart systems, such as smart home, and smart vehicle, and smart grids, critically depends on developing sophisticated planning capabilities that can anticipate the next desired function based on historical interactions. While existing methods view user behaviors as sequential data and apply models like RNNs and Transformers to predict future actions, they often fail to incorporate domain knowledge and capture personalized user preferences. In this paper, we propose a novel approach that incorporates LLM-enhanced logs and personalized prompts. Our approach first constructs a graph that captures individual behavior preferences derived from their interaction histories. This graph effectively transforms into a soft continuous prompt that precedes the sequence of user behaviors. Then our approach leverages the vast general knowledge and robust reasoning capabilities of a pretrained LLM to enrich the oversimplified and incomplete log records. By enhancing these logs semantically, our approach better understands the user’s actions and intentions, especially for those rare events in the dataset. We evaluate the method across four real-world datasets from both smart vehicle and smart home settings. The findings validate the effectiveness of our LLM-enhanced description and personalized prompt, shedding light on potential ways to advance the intelligence of smart space.
Note: While this manuscript provides description of the data, we are not permitted to make these datasets publicly available due to restrictions imposed by the data provider.
journal: AAAI
[…]
organization=MOE KLINNS Lab, Xi’an Jiaotong University, city=Xi’an, postcode=710054, country=China
[…]
abstract
1 Introduction
In today’s society, smart devices have become an indispensable part of our daily lives, significantly enriching our life experiences. However, despite the variety of functionalities these smart devices offer, they largely operate independently of one another, lacking deep inter-connectivity. This isolation hinders the formation of a truly “smart” systems. Take the smart vehicle scenario as an example: on a cold workday morning, a driver might open their car door, followed by the need to adjust the seat position, turn on the heating, set the navigation to plan a congestion-free route, and play their favorite music. While these functions are straightforward, manually executing each one can be tedious. Enabling a simple action, like opening a door, to automatically trigger all related tasks would significantly enhance the user experience in smart settings. Simplifying the interaction between users and smart devices to allow the devices to understand and automatically adapt to users’ needs in various scenarios represents an urgent yet challenging goal [1].
This vision intersects with two main areas of research: human activity prediction and sequential recommendation. In human activity prediction, previous works in the context of smart space attempt to build models from historical interaction data between users and their environments, with the goal of automating the recognition of users’ intentions and plan the next move [2, 3, 4, 5]. For instance, deep learning models have been employed to analyze various sensor data within smart homes to forecast subsequent user activities [6]. On the other hand, sequential recommendation focuses on the online activity context suggesting content that users may find engaging based on their behaviors [7, 8]. This includes analyzing sequences of user online interactions—such as movie watching, web browsing, clicking, bookmarking, and purchasing—to tailor recommendations of potentially interesting content like movies or products.
However, in real-world smart spaces, existing methods face significant challenges that differ markedly from those in recommendation systems. While recommendation systems can access extensive side information—including the names, brands, and attributes of purchased items [8], and content of movies that users are interested in—data logs from devices and sensors in smart spaces are often sparse or even incomplete. Consequently, earlier methods heavily often employ one-hot encoding to represent different events as unrelated IDs, missing the inherent connections between related actions. For example, opening a refrigerator and preheating an oven are both linked to the intention of eating, yet are treated as unrelated by one-hot encoding. This approach fails to recognize the natural associations between activities, leading to a disconnect in representations, particularly for low-frequency actions that are poorly represented.
Additionally, due to the limited number of users in previous smart space datasets [2, 9, 10, 5], prior research in smart spaces typically employs a single model to predict a range of user actions, neglecting the individual preferences that result in behavioral differences. For instance, the routine of going to sleep could involve actions like drawing the curtains and switching off the lights for some, while for others, it might include playing soft music and setting an alarm. This variation highlights the inadequacy of a one-size-fits-all approach in predicting user actions without considering individual preferences, which is crucial for effective smart space automation.
To address the challenges in predicting user behavior within smart systems, we propose a novel approach that incorporates LLM-enhanced logs and personalized prompts. This method taps into the vast general knowledge and robust reasoning capabilities of a pretrained LLM, using in-context learning to enrich the oversimplified and incomplete log records. By enhancing these logs semantically, we can better understand the user’s actions and intentions. Additionally, our approach constructs a graph that captures individual behavior preferences derived from their interaction histories. This graph effectively transforms into a soft continuous prompt that precedes the sequence of user behaviors. Through prompt tuning, the model is guided to make differentiated predictions about user behavior based on this prompt.
We primarily evaluate our approach through experiments conducted on two large-scale, real-world dataset of user interactions with smart vehicles, complemented by two additional datasets focusing on smart home settings. The experimental outcomes demonstrate the superior performance of our approach when compared to state-of-the-art methods in terms of predicting a user’s next move within smart vehicle contexts. By incorporating LLMs for description augmentation, our approach significantly boosts performance in scenarios involving low-frequency events. Furthermore, our framework demonstrates robust adaptability, effectively handling cold-start scenarios with new users and seamlessly transitioning to smart home contexts. The contributions are summarized as follows:
  • 1.
    We introduce a novel prompt-based approach for predicting user behavior, which adapts dynamically to individual user preferences, making it a pioneering application of prompt tuning for customized predictions within the smart space context.
  • 2.
    Utilizing a pre-trained LLM, our method significantly enhances event representations by augmenting them with semantic descriptions, particularly effective for addressing rare events.
  • 3.
    We validate our approach with extensive experiments on datasets from two different domains, demonstrating its effectiveness in accurately predicting user behavior.
2 Description of the Industrial Dataset
In this paper, we present two large-scale, real-world datasets from the industrial sector. These datasets comprise logs of human-vehicle interactions collected from 100 private smart automobiles in two different cities (50 vehicles per city), over a 90-day period through January to March. Each log entry captures key details such as the vehicle ID, signal name, signal value, and timestamp. In total, the datasets include over 9.55 million events generated by 357 types of sensors. These sensors monitor everyday interactive functions like gear-shift positions, with a full list provided in Table 1. To ensure privacy, all vehicle IDs have been anonymized using randomly generated strings, and the datasets do not contain any sensitive information, such as location data or navigation destinations. The collection and use of these datasets have been reviewed and approved by the local institution’s ethics review board.
Table 1: The event category in our dataset and the sensor number in each category.
Category # of Event Example
Air Conditioner 51 Heat Request
Entertainment 47 Media Volume Configuration
Energy Management 43 Fuel Level Position
Door & Window 37 Front Left Door Status
ADAS 30 APA Func Status
Remote Request 29 Remote Close Trunk Request
Seat Adjustment 19 Driver Buckle Status
Locking System 18 Tank Lock Request
Steering Wheel 16 Steer Wheel Heat Request
Signal Lighting 15 Reverse Light Status
Rear Mirror 13 Rear Mirror Fold Status
Gear & Velocity 11 Gear Shift Position
Tire Monitoring 9 Front Left Tire Pressure
In-car Lighting 7 Ambient Light Status
Window Wiper 3 Window Wiper Status
Misc. 9 Limp Home Status
The sensors capture both continuous and discrete events. Our goal is to convert these events into dense embeddings using one-hot vectors. However, creating an embedding for each value of a continuous event isn’t practical. That approach would lead to an overwhelming number of parameters, complicating and slowing down the learning process. To avoid this issue, we group similar continuous event values into groups, or ”bins”. We choose the number of bins in a way that keeps values within the same bin as close as possible to each other, while making sure that values from different bins are as distinct as possible. This discretization process transforms continuous events into a manageable set of discrete states. As a result, we can effectively learn embeddings for each event-state pair.
We also observed that there exist events that almost always co-occur. An instance of this is the co-occurrence of the reverse gear status event and the reverse light status event. Due to the function logic, these two types of events should be occur at the same time. However, since the combination of such events does not offer any additional information over a single event, we replace the combination with a single event (e.g., removing all the reverse light status events) to eliminate redundancy.
After discretizing the continuous signals, each record in our dataset includes an event’s name along with its discrete state. The state of an event can significantly influence its meaning; the same event can imply different actions or intentions based on its status. For instance, “AntiTheft_1” might suggest that the driver has exited the vehicle, whereas “AntiTheft_0” typically indicates the driver is entering the vehicle. Given this variability, it’s crucial to represent events differently depending on their status.
To achieve this, we create a unique pairing of each event with its corresponding status, resulting in 487 distinct event-status combinations. Each of these pairs is then converted into a one-hot vector. These vectors serve as input to a standard PyTorch Embedding layer, which transforms them into dense, continuous embeddings. Through training, our approach refines these embeddings so that those representing similar intentions are closer together in the embedding space. For example, “AntiTheft_0” would be nearer to “DoorLock_0” but further from “AntiTheft_1”.
In our dataset, the most frequently occurring events align closely with typical daily vehicle usage. These include events such as starting and stopping the car, shifting gears, opening and closing doors, activating the heated steering wheel, applying the brakes, adjusting windows, and toggling the seatbelt. Also, the event distribution across time of day and days of the week is displayed in Fig. 1. However, we are not permitted to make these datasets publicly available due to restrictions imposed by the data provider.
Figure 1: Event Distribution Across Time of Day and Days of the Week
3 Conclusion
In this paper, we propose a novel approach that incorporates LLM-enhanced logs and personalized prompts. Our approach constructs a graph that captures individual behavior preferences derived from their interaction histories, which effectively transforms into a soft continuous prompt to precede the sequence of user behaviors. our approach then leverages the vast general knowledge and robust reasoning capabilities of a pretrained LLM to enrich the oversimplified and incomplete log records. By enhancing these logs semantically, our approach builds more representative embeddings for events, especially for those rare ones in the dataset. Our methodology’s effectiveness has been validated through rigorous evaluation on a large-scale dataset involving user interactions with smart cars, along with two additional datasets from smart home contexts. The results indicate substantial improvements over existing SOTA methods and showcase the potential for generalization across various smart environments, demonstrating the versatile applicability and robustness of our framework in smart environment settings.